OpenStudy (anonymous):

ive tried to prove it many times but , i still didnt get it.... if dy/dx = -y/x+2y prove that d''y/dx''=8/(x+2y)^3

OpenStudy (tkhunny):

Do you mean \(\dfrac{-y}{x+2y}\)? That is NOT what you have written. Folks in calculus should know about the Order of Operations. Have you considered the Quotient Rule for differentiation?

OpenStudy (precal):

use the quotient rule

OpenStudy (anonymous):

really ? sorry i dont know how to write it in the right way...yes i use that quotient rule , but then i got this\[\frac{ -\frac{ dy }{ dx }.x+2y-(2y.\frac{ dy }{ dx }) }{ (x+2y)^2 }\] then what should i do ?

OpenStudy (tkhunny):

You should try that again. The whole numerator is a mess.

OpenStudy (anonymous):

ok..wait

OpenStudy (anonymous):

someby pls help me..i dont get it..... :{

OpenStudy (tkhunny):

Numerator: (x+2y)(-y') - (-y)(1+2y') Look at it very carefully.

OpenStudy (anonymous):

oh yeahh my numerator is really a mess..then what should i do ?

OpenStudy (anonymous):

what should i do ~~~~~ omgg

hartnn (hartnn):

even i tried this more than once, and i don't get numerator = 8, is anything else given ?

OpenStudy (anonymous):

the original question .

OpenStudy (anonymous):

idk what im doing -.-

OpenStudy (anonymous):

first question

hartnn (hartnn):

ohhh...so, xy+y^2 =4.... then its easy, did you get that numerator first ? next thing will be to simplify it.

hartnn (hartnn):

got this : (x+2y)(-y') - (-y)(1+2y') first ?

OpenStudy (anonymous):

yes i got it just now .haha

OpenStudy (anonymous):

then ?

hartnn (hartnn):

now use, y'=-y/(x+2y)

OpenStudy (anonymous):

yes im using that..then quotient rule ?

hartnn (hartnn):

(x+2y)(-y') - (-y)(1+2y') is the result of numerator AFTER applying quotient rule, now you have to simplify

OpenStudy (tkhunny):

Just for the record, it is ALWAYS more beneficial to show the ORIGINAL problem statement.

hartnn (hartnn):

yeah, that would have saved me few minutes...

OpenStudy (anonymous):

to be honest im stucking again

hartnn (hartnn):

what u got after substituting y'=-y/(x+2y) in that numerator ?

OpenStudy (anonymous):

\[2y-\frac{ 2y^2 }{ x+2y }\] what about this ? for that numerator

OpenStudy (anonymous):

\[\frac{ 2xy+2y }{ (x+2y)^3 }\]

OpenStudy (anonymous):

this ???????

OpenStudy (anonymous):

then subtitute (0,2) into that equation ?

hartnn (hartnn):

its actually this : \(\huge \frac{ 2xy+2y^2 }{ (x+2y)^3 }=\frac{ 2(xy+y^2) }{ (x+2y)^3 }\) now use xy+y^2 =4 and you are done

hartnn (hartnn):

\(2y-\frac{ 2y^2 }{ x+2y }\) was correct

OpenStudy (anonymous):

\[\frac{ 2xy+2y^2}{(x+2y)^3}\]

OpenStudy (anonymous):

this ????am i right ??

hartnn (hartnn):

yeah, factoring out 2 will give you 2 (xy+y^2) and xy+y^2 as per the question is 4

OpenStudy (anonymous):

ohhhh........i see ur answer........and i totally got it ! thanks a lot @hartnn !!:) really appreciate ^_^

hartnn (hartnn):

welcome ^_^